# CLQCD Collaboration

### v2

@@ I Introduction
Lattice QCD provides the primary non-perturbative framework for ab initio calculations of strong interaction physics. A central challenge in the field is the choice of fermion discretization, which must balance computational cost against the preservation of fundamental symmetries, most notably chiral symmetry. In this context, the mixed-action approach has emerged as a powerful pragmatic strategy. By generating gauge ensembles with a computationally efficient “sea” fermion action and calculating observables using a more theoretically clean but expensive “valence” fermion action, one can leverage the statistical advantages of large ensembles while maintaining good control over systematic errors. Mixed-action lattice QCD has been employed in calculations of glue helicity Yang:2017, nucleon charges (e.g.,Chang:2018; Jang:2020; Park:2025nucleon), hadron vacuum polarization Borsanyi:2021, and other observables.
@@
Recent systematic study Zhao:2022ooq has observed that when the sea fermion action preserves chiral symmetry, $\Delta_{\text{mix}}$ scales as $\mathcal{O}(a^{4})$—a significant improvement over the expected $\mathcal{O}(a^{2})$ behavior. This favorable scaling has been observed over a wide range of lattice spacings, $a\in[0.04,0.19]$ fm, for chirally symmetric actions such as Domain wall Brower:2004xi; Brower:2005qw; Brower:2012vk and Highly Improved Staggered Quark (HISQ) fermions Follana:2006rc. In contrast, when a clover sea fermion action with explicit chiral symmetry breaking is used, $\Delta_{\text{mix}}$ remains substantially larger, independent of the valence action.
However, the comparison in Ref. Zhao:2022ooq presents potential loopholes. The clover fermion ensembles used the tadpole-improved tree-level Symanzik (TITLS) gauge action, while the HISQ and Domain wall ensembles used different gauge actions (one-loop improved Symanzik and Iwasaki, respectively). Consequently, the possibility that the gauge action itself significantly impacts $\Delta_{\text{mix}}$ cannot be excluded. Furthermore, the number of dynamical fermion flavors differs between the clover ($N_{f}=2+1$) and HISQ ($N_{f}=2+1+1$) ensembles, introducing another confounding factor.
@@ II Setup and basic tests of the generated gauge ensembles
The fermion action used for the gauge ensembles generated in this work is the HISQ action Follana:2006rc, which effectively mitigates taste symmetry breaking and discretization errors while retaining acceptable chiral symmetry. It is formulated in the following form:
@@
  • $C_{R}=-1/(20u_{0}^{2})$ for the tadpole-improved Symanzik action, with $u_{0}=\left(\langle\mathrm{Re}\mathrm{Tr}(P)\rangle/3\right)^{1/4}$. This was employed in previous CLQCD ensembles with clover fermions CLQCD:2023sdb; CLQCD:2024yyn;
@@
In this work, we generated a set of 2+1+1-flavor gauge ensembles using the aforementioned fermion and gauge actions. The ensemble parameters are detailed in Table 1 (with the symbol $\tilde{X}\equiv Xa^{n}$ for any quantity $X$ with mass dimension $n$), including the gauge coupling $\hat{\beta}$, the tadpole factor $u_{0}$, lattice volume $\tilde{L}^{3}\times\tilde{T}$, and the bare dimensionless quark masses $\tilde{m}^{\rm b}_{l,s,c}$. For all ensembles at $\hat{\beta}=7.29$ and 7.54 (except c24P31s to be discussed later), we fixed the charm-to-strange mass ratio to $\tilde{m}^{\rm b}_{c}/\tilde{m}^{\rm b}_{s}=11.766$, corresponding to the 2+1+1 flavor FLAG average FlavourLatticeAveragingGroupFLAG:2024oxs; EuropeanTwistedMass:2014osg; Chakraborty:2014aca; FermilabLattice:2018est; ExtendedTwistedMass:2021gbo. For the ensembles at $\beta=7.75$ and 8.20, we used $\tilde{m}^{\rm b}_{c}/\tilde{m}^{\rm b}_{s}=11.74$, a value consistent with the FLAG average of 11.766(30) within its uncertainty. The strange-to-light mass ratio, $\tilde{m}^{\rm b}_{s}/\tilde{m}^{\rm b}_{l}$, was set to 27.18 Bazavov:2017lyh for the physical point ensemble (c48P13), to 10 for c24P22 and c32P22, and to 5 for the remaining ensembles (excluding c24P31s). The bare strange quark mass $\tilde{m}^{\rm b}_{s}$ was then tuned to ensure the resulting $m_{\eta_{s}}$ mass was close to the value of $689.63(18)$ MeV Borsanyi:2020mff which corresponds to the physical strange quark mass, and the tadpole factor $u_{0}$ was tuned to ensure consistency within 0.001% between its input value and the value from thermalized configurations. Finally, to investigate the impact of the strange sea quark mass, we generated the c24P31s ensemble using the same $\hat{\beta}$ and $\tilde{m}^{\rm b}_{l,c}$ as the c24P31 ensemble but with a $\tilde{m}^{\rm b}_{s}$ enlarged by 17%.
@@
We determine the lattice spacing $a_{w_{0}}$ for each ensemble using the gradient flow scale with $w_{0}=0.1725(7)$ fm, based on the FLAG average FlavourLatticeAveragingGroupFLAG:2024oxs; Dowdall:2013rya; ExtendedTwistedMass:2021qui; Miller:2020evg; MILC:2015tqx; Borsanyi:2020mff. The results are summarized in Table 2. To perform a global fit, we adopt the empirical formula from Ref. CLQCD:2023sdb, originally used for CLQCD clover fermion ensembles. The fit function is:
@@
Here, $m_{\pi,{\rm phys}}=134.98$ MeV ParticleDataGroup:2020ssz is the physical pion mass with QED corrections highly suppressed, $m_{\eta_{s},{\rm phys}}=689.63(18)$ MeV Borsanyi:2020mff corresponds to the physical strange quark mass, and $u^{I}_{0}$ denotes the input value of $u_{0}$. The fit yields a $\chi^{2}$/d.o.f. of 1.3, and the resulting parameters are also provided in Table 2.
@@
Table 3: Input parameters $\hat{\beta}$, $u_{0}$, $\tilde{L}\times\tilde{T}$, and $\tilde{m}^{\rm b}_{l,s}$ of the 2+1 flavor gauge ensembles with HISQ fermion and different gauge actions.
$C_{R}$ $\hat{\beta}$ $\tilde{L}^{3}\times\tilde{T}$ $u_{0}$ $\tilde{m}^{\rm b}_{l}$ $\tilde{m}^{\rm b}_{s}$
-1/20 6.743 $24^{3}\times 48$ 0.872958(4) 0.01027 0.05135
@@
Table 4 lists the pseudoscalar meson masses $m_{\pi}$, $m_{\eta_{s}}$, and $m_{\eta_{c}}$, determined using the jointly fitted lattice spacing. The $m_{\eta_{s}}$ values on all the 2+1+1 flavor ensembles are consistent with the our target value of 689.63(18) MeV Borsanyi:2020mff within $2\sigma$, except for the c24P31s ensemble used to test strange mass mistuning. The $m_{\eta_{c}}$ results also reproduce the physical value to within approximately 1%.
@@
  • SC+Stad: The 2+1 flavor stout-smeared clover (SC) fermion ensembles with the same Stad gauge action from the CLQCD collaboration CLQCD:2023sdb;
  • HI+S(1): The 2+1+1 flavor HISQ fermion ensembles with a one-loop-improved Symanzik (S(1)) gauge action from the MILC collaboration MILC:2010pul; MILC:2012znn (data taken from Ref. Zhao:2022ooq).
@@
For the Symanzik gauge action, the strong coupling $\alpha_{s}$ can be estimated at leading-order perturbation theory through the tadpole factor $u_{0}$ Alford:1995hw, defined as
@@ III Result of $\Delta_{\rm mix}$
All our results of $\Delta_{\text{mix}}$ are collected in Table 6, together with a generalized mixed action effect between valence fermion actions Zhao:2022ooq,
@@
The lattice spacing dependence of $\Delta_{\text{mix}}$ on the 2+1+1 flavor HISQ ensembles is shown in Fig. 3. The SC and HC results exhibit a clear $\mathcal{O}(a^{4})$ scaling across the four lattice spacings, despite a large relative uncertainty at the finest lattice spacing ($a=0.048$ fm). The overlap results, though limited to two data points, are also consistent with this $\mathcal{O}(a^{4})$ scaling behavior.
@@
Figure 4: The mixed action effects $\Delta_{\rm mix}$ at $m_{\pi}\sim 0.30$ GeV for various valence fermion actions (OV, SC and HC) and $\bar{\Delta}_{\rm mix,uni}$ for different valence fermion actions (OV+HC, OV+SC and HC+HC) on the 2+1 flavor HI+I, HI+Stad, HI+S(0) ensembles, 2+1+1 flavor HI+Stad ensembles, and also includes the values from Ref. Zhao:2022ooq (shown as histograms in the shaded region) for comparison
. All the quantities are in unit of GeV2, and are interpolated to $a=0.11$ fm and $m_{\pi}\sim 300$ MeV to make a fair comparison.
To facilitate a comparison of $\Delta_{\text{mix}}$ across different fermion-gauge action combinations and the presence of charm sea quarks, we interpolated all results to a common lattice spacing of $a=0.11$ fm, as shown in Fig. 4.
For the 2+1 flavor cases, $\Delta_{\text{mix}}$ exhibits an obvious decrease with increasing absolute value of $|C_{P}|$, rendering its value with the Iwasaki action being 0.61(4) of that with the tree level Symanzik gauge actions. Nevertheless, $\Delta_{\text{mix}}$ with Domain wall fermions is still smaller by a factor of 2.5(3) compared with HISQ fermions under the same Iwasaki gauge action. When using the same Stad gauge action, $\Delta_{\text{mix}}$ with HISQ fermions is significantly smaller than with stout clover (SC) fermions, confirming the importance of using a chirally symmetric fermion action for generating gauge configurations on suppression $\Delta_{\text{mix}}$.
@@
A comparison between the 2+1 and 2+1+1 flavor HISQ ensembles with the same Stad gauge action reveals that $\Delta_{\text{mix}}$ is consistent within statistical uncertainties, regardless of the inclusion of charm sea quarks, once the lattice spacing is rescaled to a common value assuming $a^{4}$ scaling. In contrast, both of these $\Delta_{\text{mix}}$ values are lower than that of the 2+1+1 flavor HISQ ensemble with the S(1) gauge action.

### v3

@@ I Introduction
Lattice QCD provides the primary non-perturbative framework for ab initio calculations of strong interaction physics. A central challenge in the field is the choice of fermion discretization, which must balance computational cost against the preservation of fundamental symmetries, most notably chiral symmetry. In this context, the mixed-action approach has emerged as a powerful pragmatic strategy. By generating gauge ensembles with a computationally efficient “sea” fermion action and calculating observables using a more theoretically clean but expensive “valence” fermion action, one can leverage the statistical advantages of large ensembles while maintaining good control over systematic errors. Mixed-action lattice QCD has been employed in calculations of glue helicity Yang and others (2017), nucleon charges (e.g.,Chang and others (2018); Jang and others (2020); Park and others (2025)), hadron vacuum polarization Borsanyi and others (2021a), and other observables.
@@
In the past two decades, numerical lattice QCD studies of mixed-action effects have remained relatively sparse and fragmented Durr et al. (2007); Orginos and Walker-Loud (2008); Aubin et al. (2008); Walker-Loud and others (2009); Li and others (2010); Lujan et al. (2012); Cichy et al. (2013); Basak et al. (2015); Berkowitz and others (2017); Basak et al. (2017). This is largely because such effects are typically treated as a component of the systematic uncertainty for specific valence-sea fermion combinations rather than as a primary subject of study. A systematic numerical investigation into the impact of specific valence and sea actions was absent until Ref. Zhao et al. (2023).
Based on high-precision calculations of $\Delta_{\text{mix}}$, Ref. Zhao et al. (2023) established that when the sea fermion action preserves chiral symmetry, $\Delta_{\text{mix}}$ scales as $\mathcal{O}(a^{4})$—a significant improvement over the expected $\mathcal{O}(a^{2})$ behavior. This favorable scaling has been observed over a wide range of lattice spacings, $a\in[0.04,0.19]$ fm, for chirally symmetric actions such as DW Brower et al. (2005, 2006, 2017) and HISQ fermions Follana et al. (2007). In contrast, when a clover sea fermion action with explicit chiral symmetry breaking is used, $\Delta_{\text{mix}}$ remains substantially larger, independent of the valence action.
However, the comparison in Ref. Zhao et al. (2023) presents potential loopholes. The clover fermion ensembles used the tadpole-improved tree-level Symanzik (TITLS) gauge action, while the HISQ and DW ensembles used different gauge actions (one-loop improved Symanzik and Iwasaki, respectively). Consequently, the possibility that the gauge action itself significantly impacts $\Delta_{\text{mix}}$ cannot be excluded. Furthermore, the number of dynamical fermion flavors differs between the clover ($N_{f}=2+1$) and HISQ ($N_{f}=2+1+1$) ensembles, introducing another confounding factor.
@@ II Setup and basic tests of the generated gauge ensembles
The fermion action used for the gauge ensembles generated in this work is the HISQ action Follana et al. (2007), which effectively mitigates taste symmetry breaking and discretization errors while retaining acceptable chiral symmetry. It is formulated in the following form:
@@
  • $C_{R}=-1/(20u_{0}^{2})$ for the tadpole-improved Symanzik action, with $u_{0}=\left(\langle\mathrm{Re}\mathrm{Tr}(P)\rangle/3\right)^{1/4}$. This was employed in previous CLQCD ensembles with clover fermions Hu and others (2024); Du and others (2025b);
@@
In this work, we generated a set of 2+1+1-flavor gauge ensembles using the aforementioned fermion and gauge actions. The ensemble parameters are detailed in Table 1 (with the symbol $\tilde{X}\equiv Xa^{n}$ for any quantity $X$ with mass dimension $n$), including the gauge coupling $\hat{\beta}$, the tadpole factor $u_{0}$, lattice volume $\tilde{L}^{3}\times\tilde{T}$, and the bare dimensionless quark masses $\tilde{m}^{\rm b}_{l,s,c}$. For all ensembles at $\hat{\beta}=7.29$ and 7.54 (except c24P31s to be discussed later), we fixed the charm-to-strange mass ratio to $\tilde{m}^{\rm b}_{c}/\tilde{m}^{\rm b}_{s}=11.766$, corresponding to the 2+1+1 flavor FLAG average Aoki and others (2024); Carrasco and others (2014); Chakraborty et al. (2015); Bazavov and others (2018b); Alexandrou and others (2021a). For the ensembles at $\beta=7.75$ and 8.20, we used $\tilde{m}^{\rm b}_{c}/\tilde{m}^{\rm b}_{s}=11.74$, a value consistent with the FLAG average of 11.766(30) within its uncertainty. The strange-to-light mass ratio, $\tilde{m}^{\rm b}_{s}/\tilde{m}^{\rm b}_{l}$, was set to 27.18 Bazavov and others (2018a) for the physical point ensemble (c48P13), to 10 for c24P22 and c32P22, and to 5 for the remaining ensembles (excluding c24P31s). The bare strange quark mass $\tilde{m}^{\rm b}_{s}$ was then tuned to ensure the resulting $m_{\eta_{s}}$ mass was close to the value of $689.63(18)$ MeV Borsanyi and others (2021b) which corresponds to the physical strange quark mass, and the tadpole factor $u_{0}$ was tuned to ensure consistency within 0.001% between its input value and the value from thermalized configurations. Finally, to investigate the impact of the strange sea quark mass, we generated the c24P31s ensemble using the same $\hat{\beta}$ and $\tilde{m}^{\rm b}_{l,c}$ as the c24P31 ensemble but with a $\tilde{m}^{\rm b}_{s}$ enlarged by 17%.
@@
We determine the lattice spacing $a_{w_{0}}$ for each ensemble using the gradient flow scale with $w_{0}=0.1725(7)$ fm, based on the FLAG average Aoki and others (2024); Dowdall et al. (2013); Alexandrou and others (2021b); Miller and others (2021); Bazavov and others (2016); Borsanyi and others (2021b). The results are summarized in Table 2. To perform a global fit, we adopt the empirical formula from Ref. Hu and others (2024), originally used for CLQCD clover fermion ensembles. The fit function is:
@@
Here, $m_{\pi,{\rm phys}}=134.98$ MeV Zyla and others (2020) is the physical pion mass with QED corrections highly suppressed, $m_{\eta_{s},{\rm phys}}=689.63(18)$ MeV Borsanyi and others (2021b) corresponds to the physical strange quark mass, and $u^{I}_{0}$ denotes the input value of $u_{0}$. The fit yields a $\chi^{2}$/d.o.f. of 1.3, and the resulting parameters are also provided in Table 2.
@@
Table 3: Input parameters $\hat{\beta}$, $u_{0}$, and $\tilde{m}^{\rm b}_{l,s}$ of the 2+1 flavor gauge ensembles with size $\tilde{L}\times\tilde{T}=24^{3}\times 48$ HISQ fermion and different gauge actions.
Ensemble $C_{R}$ $\hat{\beta}$ $u_{0}$ $\tilde{m}^{\rm b}_{l}$ $\tilde{m}^{\rm b}_{s}$
x24P31 -1/20 6.743 0.872958(4) 0.01027 0.05135
@@
Table 4 lists the pseudoscalar meson masses $m_{\pi}$, $m_{\eta_{s}}$, and $m_{\eta_{c}}$, determined using the jointly fitted lattice spacing. The $m_{\eta_{s}}$ values on all the 2+1+1 flavor ensembles are consistent with the our target value of 689.63(18) MeV Borsanyi and others (2021b) within $2\sigma$, except for the c24P31s ensemble used to test strange mass mistuning. The $m_{\eta_{c}}$ results also reproduce the physical value to within approximately 1%.
@@
  • SC+Stad: The 2+1 flavor stout-smeared clover (SC) fermion ensembles with the same Stad gauge action from the CLQCD collaboration Hu and others (2024);
  • HI+S(1): The 2+1+1 flavor HISQ fermion ensembles with a one-loop-improved Symanzik (S(1)) gauge action from the MILC collaboration Bazavov and others (2010, 2013) (data taken from Ref. Zhao et al. (2023)).
@@
For the Symanzik gauge action, the strong coupling $\alpha_{s}$ can be estimated at leading-order perturbation theory through the tadpole factor $u_{0}$ Alford et al. (1995), defined as
@@ III Result of $\Delta_{\rm mix}$
All our $2+1+1$ flavor results of $\Delta_{\text{mix}}$ are collected in Table 6, together with a generalized mixed action effect between valence fermion actions Zhao et al. (2023),
@@
The lattice spacing dependence of $\Delta_{\text{mix}}$ on the 2+1+1 flavor HISQ ensembles is shown in Fig. 3. The solid (dashed) line passes through the HC/HI $\Delta_{\mathrm{mix}}$ point at the coarsest spacing and indicates $a^{4}$ ($a^{2}$) scaling. The SC and HC results exhibit a clear $\mathcal{O}(a^{4})$ scaling across the four lattice spacings, despite a large relative uncertainty at the finest lattice spacing ($a=0.048$ fm). The overlap results, though limited to two data points, are also consistent with this $\mathcal{O}(a^{4})$ scaling behavior. We further parameterize $\Delta_{\text{mix}}$ with the following form,
$\Delta_{\text{mix}}(a)=c_{2}a^{2}+c_{4}a^{4},$ (9)
and list the fitted (or solved in the overlap case) coefficients $c_{2,4}$ in Table 6. It is evident that the values of $c_{2}$ are significantly non-zero in cases involving the stout-link smeared clover action (SC), although they remain much smaller than the corresponding $c_{4}$ values. Since $a^{4}$ scaling is observed for both the exactly chiral overlap fermion and the HYP-smeared clover fermion (with its suppressed chiral symmetry breaking), the scaling can be strongly related to chiral symmetry.
@@
Figure 4: The mixed action effects $\Delta_{\rm mix}$ at $m_{\pi}\sim 0.30$ GeV for various valence fermion actions (OV, SC and HC) and $\bar{\Delta}_{\rm mix,uni}$ for different valence fermion actions (OV+HC, OV+SC and HC+HC) on the 2+1 flavor HI+I, HI+Stad, HI+S(0) ensembles, 2+1+1 flavor HI+Stad ensembles, and also includes the values from Ref. Zhao et al. (2023) (shown as histograms in the shaded region) for comparison. All the quantities are in unit of GeV2, and are interpolated to $a=0.112$ fm and $m_{\pi}\sim 300$ MeV to make a fair comparison.
Given that the lattice spacings of the x/z24P31 ensembles (aa=0.111/0.113 fm) differ from that of the y24P31 ensemble (aa=0.112 fm) by only 1%, we interpolated all the $2+1+1$ flavor results to a common reference point of $a=0.112$ fm using the parameterization in Eq. (9). This enables a direct comparison of $\Delta_{\text{mix}}$ across different fermion-gauge action combinations and charm quark content. The interpolated results are presented in Fig. 4.
@@
A comparison between the 2+1 and 2+1+1 flavor HISQ ensembles with the same Stad gauge action reveals that $\Delta_{\text{mix}}$ is consistent within statistical uncertainties, regardless of the inclusion of charm sea quarks. In contrast, both of these $\Delta_{\text{mix}}$ values are lower than the values Zhao et al. (2023) on the 2+1+1 flavor HISQ ensemble with the S(1) gauge action.
For the 2+1 flavor cases, $\Delta_{\text{mix}}$ exhibits an obvious decrease with increasing absolute value of $|C_{P}|$, rendering its value with the Iwasaki action being 0.61(4) of that with the tree level Symanzik gauge actions. Nevertheless, $\Delta_{\text{mix}}$ with DW fermions is still smaller by a factor of 2.5(3) compared with HISQ fermions under the same Iwasaki gauge action. When using the same Stad gauge action, $\Delta_{\text{mix}}$ with HISQ fermions is significantly smaller than with stout clover (SC) fermions, confirming the importance of using a chirally symmetric fermion action for generating gauge configurations on suppression $\Delta_{\text{mix}}$.
The preceding calculations enable us to isolate the specific origins of variations in $\Delta_{\text{mix}}$ by comparing our results with existing literature:
[…]
3. OV@HI2+1+1+S(1) Basak et al. (2015): Using a 2+1+1 flavor HISQ ensemble with a 1-loop Symanzik gauge action at $m_{\pi}=310$ MeV and $a\simeq 0.12$ fm, $\Delta_{\text{mix}}$ for a 1-step HYP-smeared overlap valence fermion is 0.031(3) GeV2 at $a=0.112$ fm, based on the $a^{4}$ scaling verified for this setup in Ref. Zhao et al. (2023) (note: taste-breaking effects were unsubtracted Basak et al. (2017)). This value is in excellent agreement with the 0.027(3) GeV2 from Ref. Zhao et al. (2023).
4. DW@HI2+1+1+S(1) Berkowitz and others (2017): Using 2+1+1 flavor HISQ ensembles with a 1-loop Symanzik gauge action at $m_{\pi}\in[220,310]$ MeV and $a\in[0.09,0.15]$ fm, $\Delta_{\text{mix}}$ for a Mobius domain wall valence fermion is 0.024(1) GeV2 at $a=0.112$ fm, based on an $a^{2}+a^{4}$ fit Eq. (9) with coefficients $c_{2}=0.04(1)$ GeV4 and $c_{4}=0.09(3)$ GeV6. This is slightly smaller than the OV@HI2+1+1+S(1) value, suggesting valence Mobius domain wall fermions may be a better choice on HISQ ensembles at coarser lattice spacings. However, the significantly non-zero $c_{1}$ term warrants further investigation.
5. DW@AS2+1+S(1) Orginos and Walker-Loud (2008); Aubin et al. (2008); Walker-Loud and others (2009): Using 2+1 flavor Asqtad-improved staggered (AS) sea quarks with a 1-loop Symanzik gauge action at $m_{\pi}\in[260,510]$ MeV and $a\in[0.09,0.13]$ fm, $\Delta_{\text{mix}}$ for a domain wall valence fermion is 0.065(5) GeV2 at $a=0.112$ fm. The data were fit with an $a^{2}+a^{4}$ ansatz, but the uncertainty was too large to distinguish between the scaling terms. This value is roughly twice that of the most similar case (OV@HI2+1+1+S(1)), a difference we attribute to the AS sea action, given that domain wall and overlap fermions yield similar $\Delta_{\text{mix}}$ on HISQ ensembles.
6. OV@DW2+1+I Li and others (2010); Lujan et al. (2012): Using 2+1 flavor domain wall sea quarks with an Iwasaki gauge action at $m_{\pi}\in[300,400]$ MeV and $a\in[0.082,0.114]$ fm, $\Delta_{\text{mix}}$ for a 1-step HYP-smeared overlap valence fermion is 0.010(2) GeV2 at $a=0.112$ fm. This is larger than the 0.0053(4) GeV2 reported in Ref. Zhao et al. (2023), as the partially quenched chiral extrapolation used in Ref. Lujan et al. (2012) can about twice inflate the value.
7. BC@AS2+1+S(1) Basak et al. (2017): Using 2+1 flavor AS sea quarks with a 1-loop Symanzik gauge action at $m_{\pi}\sim 300$ MeV, $\Delta_{\text{mix}}$ for a Borici-Creutz valence fermion Creutz (2008); Borici (2008) was 0.017(6) GeV2 at $a=0.15$ fm and consistent with zero (0.02(5) GeV2) at $a=0.13$ fm. Considering the tendency of the AS sea action and S(1) gauge action to enlarge $\Delta_{\text{mix}}$, the Borici-Creutz fermion appears to be a promising candidate for mixed-action setups, though higher-precision calculations at finer lattice spacings are required for confirmation.